Solve the linear system by using substitution method in -5x+3y=51 and y=10x-8
Answer by kennyxx12(3)

ok one is already isolated for y
y=10x-8
substitude it y=10x-8 this equals to y
now plug it in
-5x+3y=51(original formulae)
-5x+3(10x-8)=51(after substitution)
-5x+30x-24=51
25x-24=51
ADD 24 TO BOTH SIDES
25X=75
DIVIDE
X=3
NOW SUBSTITUTE X ON THE OTHER FORMULAE
Y=10(3)-8
Y=30-8
Y=22
X=3
ANSWER (3,22)